OpenStudy (anonymous):

For the following reaction, identify the element that was oxidized, the element that was reduced, and the reducing agent. Give an explanation for each answer. 2CO + 2NO yields 2CO2 + N2

OpenStudy (aaronq):

First, you have to assign oxidation numbers (also called states) to all the atoms in the equation and see which change (in oxidation number) from the reactants side to the products side.

OpenStudy (anonymous):

? @aaronq

OpenStudy (aaronq):

This is how you do it. You start off with a basic set of rules and you figure the rest out with simple arithmetic. Rules (most often true, except in rare occasions): Group 1 always +1 Group 2 always +2 Oxygen always -2 Group 7 always -1 Noble gases, elemental or homonuclear (the same element) molecules are zero The rest you have to figure out by subtracting or adding to achieve whatever charge the compound bears. eg. H2O H is group 1, so H = +1, there are 2 of them O must balance the charge so, O=-2 HCl H=+1, Cl=-1 MgSO4, you have to know that SO4^2- is a polyatomic ion, treat that as it's own separate thing: O=-2 Total contribution of O=-2(4) = -8 ^4 oxygen atoms Charge on ion=-2, so S has to fulfil the rest which is 6, so S=+6 back to MgSO4, so we treat SO4 as entity and it bears a -2 charge, Mg therefore has to be +2 to balance the charges. Try some: NH4+ NO2 N2O PCL3 CuBr2

OpenStudy (aaronq):

Give it a try and I'll help you correct any mistakes you make